OpenStudy (anonymous):

-8+8(-4m+3)=272 solve for m

OpenStudy (michele_laino):

we have to apply the distributive property of multiplication over addition first

OpenStudy (deadly_roses):

should be m= -8

OpenStudy (michele_laino):

hint: \(8(-4m+3) = 8 \cdot(-4)m+ 8\cdot 3=...\)